OpenStudy (anonymous):

Find the sum of the geometric series by using formula. 1-3+9-27+81-243+729-2187

OpenStudy (anonymous):

@jim_thompson5910 can you help

jimthompson5910 (jim_thompson5910):

what's the first term

OpenStudy (anonymous):

the first term is 1

jimthompson5910 (jim_thompson5910):

what's the common ratio

OpenStudy (anonymous):

Common ratio is -3

jimthompson5910 (jim_thompson5910):

how many terms are we summing

jimthompson5910 (jim_thompson5910):

or adding up

OpenStudy (anonymous):

8

jimthompson5910 (jim_thompson5910):

so a = 1 r = -3 n = 8

jimthompson5910 (jim_thompson5910):

Sn = a*(1-r^n)/(1-r) S8 = 1*(1-(-3)^8)/(1-(-3)) S8 = 1*(1-(-3)^8)/(4) S8 = 1*(1-6561)/(4) S8 = 1*(-6560)/(4) S8 = (-6560)/(4) S8 = -1640

jimthompson5910 (jim_thompson5910):

S8 = sum of first 8 terms so the sum of the first 8 terms is -1640 sure enough, you can add up 1-3+9-27+81-243+729-2187 to get -1640 and that confirms the answer

OpenStudy (anonymous):

okay I got it. Not sure why i thought n was 9. so I was getting the wrong #.

jimthompson5910 (jim_thompson5910):

ah i see

OpenStudy (anonymous):

thank you!!

jimthompson5910 (jim_thompson5910):

you're welcome